what is the method to solve 2x-y=8 and 7x-3y=28 the system of linear equation
Answer by mananth(16946)

2 x - y = 8 .............1
Value of tickets
7 x -3 y = 28 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 1
6 x -3 y = 24
7 x -3 y = 28
Add the two equations
13 x = 52
/ 13
x = 4
plug value of x in (1)
2 x - y = 8
8 - y = 8
- y = 8 -8
- y = 0
y = 0
x= 4
y= 0
m.ananth@hotmail.ca